# MAC: A Benchmark for Multiple Attribute Compositional Zero-Shot Learning

### v2

@@ 6 Discussion and Conclusion
In this paper, we address the limitations of the existing compositional zero-shot learning (CZSL) dataset and introduce the Multi-Attribute Composition (MAC) dataset, which focuses on multi-attribute composition prediction. Unlike existing datasets with single and incomplete attribute annotations, MAC offers comprehensive and representative attribute annotations. It contains a large number of compositions with multiple attributes, allowing for robust evaluation of compositional learning. We evaluate nine existing methods on MAC and propose MVP-Integrator as a strong baseline. Our work contributes to the advancement of compositional learning by highlighting the significance of multi-attribute compositions in real-world scenarios. The MAC dataset provides a valuable resource for further research and development in compositional learning-related tasks. Future work could explore the fine-grained correspondence between images and multiple attributes.

### v4

@@ VI Conclusion
In this paper, we address the limitations of the existing compositional zero-shot learning (CZSL) dataset and introduce the Multi-Attribute Composition (MAC) dataset, which focuses on multi-attribute composition prediction. Unlike existing datasets with single attribute annotations, MAC offers comprehensive and representative attribute annotations. It contains a large number of compositions with multiple attributes, allowing for robust evaluation of compositional learning. We evaluate nine methods on MAC and propose MVP-Integrator as a strong baseline. Our work contributes to the compositional learning by highlighting the significance of multi-attribute compositions in real-world scenarios. MAC dataset provides a valuable resource for further research in compositional learning tasks. Future work could explore the fine-grained correspondence between images and multiple attributes.